OpenStudy (anonymous):

prove identity. 4cot theta/1+cot^2 theta = 2sin2theta

OpenStudy (anonymous):

\[\frac{4 \cot (\theta)}{cosec^2 (\theta)} \implies 4 \frac{\cos(\theta)}{\sin (\theta)} \times \sin^2 (\theta) = 2 \times 2\sin(\theta)\cos(\theta) = 2\sin(2\theta)\]

OpenStudy (anonymous):

Using: \[\cot(\theta) = \frac{\cos(\theta)}{\sin(\theta)}\] \[\frac{1}{cosec^2(\theta)} = \sin^2 (\theta)\]

OpenStudy (anonymous):

\[(4\cos \theta/\sin \theta)/cosec^2 \theta\] =\[4\cos \theta \sin \theta \sin \theta / \sin \theta\] =\[2 \times 2\sin \theta \cos \theta = \sin 2 \theta\]

OpenStudy (anonymous):

damn u r fast

OpenStudy (anonymous):

thank you all ;]